Suppose you are given the following data:

            Currency in circulation                                                                                                                                    $1,000

            Net Demand deposits at chartered banks                                                                                                      $13,000

            Nonpersonal notice and personal savings deposits at chartered banks   $12,000

            Deposits at other financial institutions $6,000

            Nonpersonal term and foreign currency deposits                                                                                           $9,000

What are M1, M2, M2+, and M3? Assume adjustments to the various aggregates are zero.
Answer #1

M1 = currency in circulation + demand deposits

M1 = 1000 + 13000 = 14000

M2 = M1 + Nonpersonal notice and personal savings deposits at chartered bank

M2 = 14000 + 12000 = 26000

M2+ = M2 +deposits at other financial institutions

M2+ = 26000 + 6000 = 32000

M3 = M2+ +Non personal term and foreign currency deposits

M3 = 32000 + 9000 = 41000